OpenStudy (anonymous):

The interior angles of a square measure __________ degrees. 180 45 90 360

OpenStudy (anonymous):

What angles are in a square

OpenStudy (anonymous):

triangle?

OpenStudy (jdoe0001):

check with a protractor or check at http://www.mathsisfun.com/quadrilaterals.html click on "square", see the measure of each interior angle

OpenStudy (anonymous):

A square has 4 right angles

OpenStudy (anonymous):

A right angle has a degree of 90, for angles in a square means 4*90

OpenStudy (anonymous):

is it 180

OpenStudy (anonymous):

90+90+90+90

OpenStudy (anonymous):

360

OpenStudy (anonymous):

so thats it

OpenStudy (anonymous):

well

OpenStudy (anonymous):

In the figure below, if segment BC = 7 cm and segment AD = 3x - 6, then x = 3. Imported Asset

OpenStudy (anonymous):

Whats the question

OpenStudy (anonymous):

In the figure below, if segment BC = 7 cm and segment AD = 3x - 6, then x = 3.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

idk what to do

OpenStudy (anonymous):

Is it a free response answer?

OpenStudy (anonymous):

i guess if its right

OpenStudy (anonymous):

From the picture the statement you posted said 7=3x - 6, where x=3

OpenStudy (anonymous):

But that is a false statement